a)Bob ltd has issued a 6year bond which was purchased by exim bank. This bond has a par value of $80,000 paying a coupon rate of 20% every two months with a yield to maturity of 17.09%. what price will exim bank pay.
b)a firm is considering to buy a machine at $80000. this machine is expected to generate net cash inflow of $28000 each year for six years, after which is expected to be sold at $5oooo.the firm expects to earn at least 14.05% per annum.
5. use NPV and internal rate of return criteria to advise the company on whether the machine should be acquired or rejected.
1
Expert's answer
2020-11-25T17:54:55-0500

a)"C=80 000\times0.2=16 000"

m=6

"m=\frac{12}{2}=6"

 find price by formula:

"PV=\frac{\frac{C}{m}}{(1+\frac{r}{m})}+\frac{\frac{C}{m}}{(1+\frac{r}{m})^2}+\frac{\frac{C}{m}}{(1+\frac{r}{m})^3}+\frac{\frac{C}{m}}{(1+\frac{r}{m})^4}+\frac{\frac{C}{m}}{(1+\frac{r}{m})^5}+\frac{\frac{C}{m}}{(1+\frac{r}{m})^6}+\frac{\N}{(1+\frac{r}{m})^6}"


"PV=\frac{\frac{16000}{6}}{(1+\frac{0.1709}{6})}+\frac{\frac{16000}{6}}{(1+\frac{0.1709}{6})^2}+\frac{\frac{16000}{6}}{(1+\frac{0.1709}{6})^3}+\frac{\frac{16000}{6}}{(1+\frac{0.1709}{6})^4}+\frac{\frac{16000}{6}}{(1+\frac{0.1709}{6})^5}+\frac{\frac{16000}{6}}{(1+\frac{0.1709}{6})^6}+\frac{80000}{(1+\frac{0.1709}{6})^6}=82112.48"

b)




NPV is found by the formula:

"NPV=\sum\frac{CFt}{(1+i)^n}-IC"

IC =-80 000

CF=28 000

i=14.05%

the sum of all discounted cash flows (inflows and outflows) associated with the investment project

discount on line 10 and on line 12 find the required amount


IRR is the interest rate at which the present value of future cash receipts and the value of the initial investment equalize, the net present value (NPV) is 0.

IRR is found by the formula:

"NPV=\sum\frac{CFt}{(1+i)^n}-IC=0"

Find IRR in line 15


NPV>0, IRR is higher than the cost of capital, then the project does not need to be abandoned